OpenStudy (anonymous):

lim as h--->0 (find in terms of constant a) sqrt(3(a+h))-sqrt(3a) / h I know I have to multiply by the conjugate, but how do I cancel the bottom?? please help!!!

OpenStudy (anonymous):

\[\frac{ \sqrt{3a+h}-\sqrt{3a} }{ h }\]

OpenStudy (anonymous):

is that what you mean?

OpenStudy (anonymous):

yes :)

OpenStudy (anonymous):

as h--->0

OpenStudy (anonymous):

ok just multiply by the conjugate like you said first so multiply by \[\sqrt{3a+h}-\sqrt{3a}\]

OpenStudy (anonymous):

\[\sqrt{3a+h}+\sqrt{3a}\]

OpenStudy (anonymous):

sorry plus 3a not minus

OpenStudy (anonymous):

wait you mean + \[\sqrt{3a}\]

OpenStudy (anonymous):

yeah thanks LG2

OpenStudy (anonymous):

np:) just go on..

OpenStudy (anonymous):

ok I did that and then I cancelled out the squares with the square roots and I was left with (3h) / \[\sqrt{3(a+h)}\] +\[\sqrt{3a}\]

OpenStudy (anonymous):

I'm not sure what to do from there :/

OpenStudy (anonymous):

you might have to retype that after I multiply by the conjugate I have \[\frac{ 3a+3h-3a }{ h \sqrt{3(a+h}+\sqrt{3a} }\]

OpenStudy (anonymous):

the 3as cancel

OpenStudy (anonymous):

yayyy I got that right :D but how do the square roots on the bottom cancel?:O

OpenStudy (anonymous):

you cancel out 1 h so you have\[\frac{ 2h }{ \sqrt{3(a+h)}+\sqrt{3a} }\]

OpenStudy (anonymous):

you just have to plug in h and then rationalize the denominator

OpenStudy (anonymous):

so you really have\[\frac{ 2h }{ \sqrt{3a}+\sqrt{3a} }\]

OpenStudy (anonymous):

or well zero on top

OpenStudy (anonymous):

so the limit is 0

OpenStudy (anonymous):

most likely more exactly the limit is\[\frac{ 0 }{ 2\sqrt{3a} }\] since we dont know what a is

OpenStudy (anonymous):

and simplifying the denominator we are left with\[\frac{ 0 }{ 6a}\]

OpenStudy (anonymous):

ohh ok so the h on the bottom became 0 right so that's why you had \[\sqrt{3a}\] + \[\sqrt{3a}\]

OpenStudy (anonymous):

wait how did you get 6a??

OpenStudy (anonymous):

yeah it became zero because we used direct substitution

OpenStudy (anonymous):

we multiplied \[2\sqrt{3a}*\sqrt{3a}\] we were left with 3a*2

OpenStudy (anonymous):

which is 6a

OpenStudy (anonymous):

ohh ok so you multiplied to get rid of the square root?

OpenStudy (anonymous):

uhm just a little correction: